OpenStudy (anonymous):

need help quick! elijah purchased a home for 123,700. the home appreciates 5.5% each year. what is the value of the home after 13 years?

OpenStudy (anonymous):

@amistre64 is the correct formula for this y=ab^x?

RhondaSommer (rhondasommer):

yes

RhondaSommer (rhondasommer):

a=price b=appreciation x=year y=total

OpenStudy (anonymous):

i didnt get a reasonable answer, im not sure what i did wrong.

OpenStudy (anonymous):

okay so 123700x.055x13?

RhondaSommer (rhondasommer):

no

RhondaSommer (rhondasommer):

\[123700((5.5)^13)\]

OpenStudy (anonymous):

oh my bad, 123700x.055^13

RhondaSommer (rhondasommer):

yeas

RhondaSommer (rhondasommer):

sorry it only put the 1 not the 3 in the exponent part

OpenStudy (anonymous):

but then i get lime 6.6194

OpenStudy (anonymous):

how is that right?

RhondaSommer (rhondasommer):

hold on

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

to increase a number by \(5.5\%\) multiply it by \(100\%+5.5\%=105.5\%=1.055\)

OpenStudy (anonymous):

i just used a different formula, this one a=pe^rt and got 131,276.

OpenStudy (anonymous):

to do it 13 times in a row, multiply it by \(1.055^{13}\)

OpenStudy (anonymous):

okay im following ^ satellite

RhondaSommer (rhondasommer):

so sorry :( thought i understood ... my bad

OpenStudy (anonymous):

that would be 2.0057

OpenStudy (anonymous):

take a calculator and use \[123700\times (1.055)^{13}\]

OpenStudy (anonymous):

248114?

OpenStudy (anonymous):

you want me to check?

OpenStudy (anonymous):

i think its right. thank you! so what is the formula a=pe^rt used for?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=123700*1.055^13

OpenStudy (anonymous):

i have another problem but using depreciation, can i create a new question and you can take a look?

OpenStudy (anonymous):

that is for continuous compounding

OpenStudy (anonymous):

it said \(5.5\%\) per year, not continuously

OpenStudy (anonymous):

okay i see. thank you! i will tag you in my next question.

OpenStudy (anonymous):

k